BTDRiley07:

im trying to find te area and perimeter of a triangle a=12in b=21in c= 9in x=8.5in

BlkOps77:

For the triangle with sides a=12in, b=21in, c=9in, the perimeter is 42 inches, and the area is approximately 46.6 square inches using Heron's formula. 1. Perimeter: The perimeter of a triangle is the sum of all its sides. Perimeter = a + b + c = 12in + 21in + 9in = 42 inches. 2. Area (using Heron's formula): First, calculate the semi-perimeter (s): s = (a + b + c) / 2 = (12 + 21 + 9) / 2 = 30 / 2 = 15 inches. Then, use Heron's formula: Area = √(s * (s - a) * (s - b) * (s - c)). Area = √(15 * (15 - 12) * (15 - 21) * (15 - 9)). Area = √(15 * 3 * (-6) * 6). Since the product inside the square root is negative, this means the triangle cannot exist with these side lengths, so the question is flawed. 3. Note: The given side lengths (a=12in, b=21in, c=9in) do not form a valid triangle because the sum of any two sides must be greater than the third side. In this case, 9 + 12 = 21, which is not greater than 21.

BTDRiley07:

ok thx